Consider the function f(x)=sin x cos x+5 on the interval (0, 2pi).

(a) find the open interval(s) on which the function increasing or decreasing, and

(b) apply the first derivative test to identify the relative extrema.



1
Expert's answer
2021-12-02T18:48:48-0500

Let us consider the function "f(x)=\sin x \cos x+5=\frac{1}{2}\sin(2x)+5" on the interval "(0, 2\pi)."


(a) Let us find the open intervals on which the function increasing or decreasing. Since "f'(x)=\cos(2x)," and "\cos(2x)=0" implies "2x=\frac{\pi}2+\pi n," that is "x=\frac{\pi}4+\frac{\pi}2 n," we conclude that "f'(x)>0" on the open intervals "(0,\frac{\pi}4),\ (\frac{3\pi}4,\frac{5\pi}4)", "(\frac{7\pi}4, 2\pi)," and "f'(x)<0" on the open intervals "(\frac{\pi}4, \frac{3\pi}4),\ (\frac{5\pi}4,\frac{7\pi}4)." It follows that the function "f(x)" increasing on the open intervals "(0,\frac{\pi}4),\ (\frac{3\pi}4,\frac{5\pi}4)", "(\frac{7\pi}4, 2\pi)," and decreasing on the open intervals "(\frac{\pi}4, \frac{3\pi}4),\ (\frac{5\pi}4,\frac{7\pi}4)."


(b) Let us identify the relative extrema. It follows that the points "\frac{\pi}4, \ \frac{5\pi}4" are the points at which the function has the local maximum, and the points "\frac{3\pi}4, \ \frac{7\pi}4" are the points at which the function has the local minimum.